OpenStudy (anonymous):

1. What is the solution of the equation? Square root of x+10 –9 = –7 Possible answers: –6 4 14 –8

OpenStudy (anonymous):

your equation is not clear sqrt(x) +10 - 9 = -7? or sqrt(x+10) - 9 = -7? or sqrt(x+10-9) = -7?

OpenStudy (anonymous):

Use brackets!!!

OpenStudy (anonymous):

Oh ok its the square root of (x+10)

OpenStudy (anonymous):

OK well first bring the 9 to the other side

OpenStudy (anonymous):

So its the square root of (x+10)=2??

OpenStudy (anonymous):

aren't ya gonna answer?

OpenStudy (anonymous):

Sorry was afk. Yes. Now square both sides.

OpenStudy (anonymous):

Got it thx!